Circular No. 7163

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999cb IN MARKARIAN 881 B. H. Granslo, Institute of Theoretical Astrophysics, Oslo, reports the

discovery by Odd Trondal, Oslo, of an apparent supernova in Mkn 881 (= MCG +7-34-28) on unfiltered CCD images taken on Apr. 4.11 (mag 16.1), Apr. 10.05 (15.7), May 1.06 (16.8), and May 6.00 UT (17.0). An independent discovery has also been reported by M. Schwartz, Cottage Grove, OR, from an unfiltered CCD frame taken on May 8.357 (confirmed on May 11.231) with the 0.35-m Tenagra I automated supernova patrol telescope in Sonoita, AZ, the new star appearing at mag ∼ 18.2. Granslo provides the following position and magnitude for SN 1999cb from CCD images taken by R. J. Irgens and H. Holen on May 9.88 with the Nordic Optical Telescope at La Palma: α = 16^{h}25^{m}48^{s}.41, δ = +40^{o}20′32′′.4 (equinox 2000.0), V = 17.9; this is 11′′ west and 10′′ south of the galaxy nucleus. Schwartz provides the following position end figures for SN 1999cb: 48^{s}.45, 32′′.3. Schwartz adds that the new object is absent from archive images taken on 1998 June 14 and July 19 using the same equipment; the limiting magnitude on his discovery, verification, and archive images is ∼ 19.

SUPERNOVA 1999cc IN NGC 6038 Schwartz also reports his discovery of an apparent supernova (mag ∼

16.3) in NGC 6038 on an unfiltered CCD frame taken on May 8.248 UT (confirmed on May 11.234) with the Tenagra I telescope. The new object is located at α = 16^{h}02^{m}42^{s}.03, δ = +37^{o}21′34′′.4 (equinox 2000.0), which is ∼ 16′′.8 east and 1′′.7 north of the galaxy nucleus. SN 1999cc was also confirmed on a frame taken May 11.087 by T. Boles, Wellingborough, U.K. The new star does not appear on either the first or the second Palomar Sky Survey scans, and it is absent from Schwartz’s archive images taken on 1998 June 15, July 20, and Sept. 14; the limiting magnitude on his discovery, verification, and archive images is ∼ 19.

V4444 SAGITTARII Visual magnitude estimates, reported in part by J. A. Mattei, AAVSO:

May 3.125 UT, 10.1 (M. Kohl, Laupen, Switzerland); 4.600, 10.2 (R. Carver, Wombarra, NSW); 5.385, 10.5 (R. Modic, Richmond Heights, OH); 6.592, 10.6 (R. Stubbings, Drouin, Victoria); 9.076, 10.9 (P. Schmeer, Bischmisbreak heim, Germany); 10.099, 11.0 (Schmeer).

1999 May 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green